OpenStudy (anonymous):

supply reasons for each of the first four steps given here. (575+37)+75=525+(37+75) =525+(75+37) =(525+75)+37 =600+37 =637

OpenStudy (anonymous):

\[(575+37)+75=525+(37+75)\] the parentheses moved that is the associative law of addition

OpenStudy (anonymous):

\[525+(37+75) =525+(75+37)\] the \(37\) and \(75\) switched places that is the commutative law of addition

OpenStudy (anonymous):

\[ 525+(75+37) =(525+75)+37\] parentheses moved again again the associative law

OpenStudy (anonymous):

\[(525+75)+37 =600+37\] is not property, that is addition as \(525+75=600\)

OpenStudy (anonymous):

as is \(600+37=637\)

OpenStudy (anonymous):

this seemed like pretty easy problem at first